OpenStudy (anonymous):

ok new problem i need to figure out f^-1(9) that is being plugged into the problem f(r)=pir^2 i know f^-1 means the inverse i just can't seem to get an answer for this problem

OpenStudy (mathstudent55):

\(f(r) = \pi r^2\) You understand that f(9) means plug in 9 for r in the formula above, and find the value of f, right?

OpenStudy (anonymous):

yes i understand that much its just the inverse part that is screwing me up we having gone over much of it yet in our class

OpenStudy (mathstudent55):

\(f^{-1}(r)\) means plug in the value of f and find r. The equation you have is the formula for the area of a circle. Just like f(9) means "what is the area of the circle with radius 9?" \(f^{-1}\) means what is the radius for a given area.

OpenStudy (anonymous):

ok i understand that but what would i get as a answer what i keep getting doesn't work

OpenStudy (freckles):

you have area=9 and you want the radius such that the area=9 That means you are to solve pi*r^2=9 for r

OpenStudy (freckles):

what do you get for r?

OpenStudy (anonymous):

this is what I'm not sure of i don't understand how to bring this further

OpenStudy (freckles):

You want to solve for r But you could solve for r^2 first. Can you solve pi*r^2=9 for r^2?

OpenStudy (freckles):

All you need is one step to do that

OpenStudy (freckles):

hint: r^2 is being multiplied by pi

OpenStudy (mathstudent55):

\(f(r) = \pi r^2\) \(9 = \pi r^2\) \(\dfrac{9}{\pi} = r^2\) \(r = \sqrt{\dfrac{\pi}{9}}\) \(r = \dfrac{\sqrt{\pi}}{3} \)

OpenStudy (freckles):

well I think @mathstudent55 gave you the recipocal of the answer

OpenStudy (anonymous):

ok well i understand it a lot more then i did before thanks again for all your help

OpenStudy (mathstudent55):

LOL. Sorry. I couldn't see as I was typing bec OS is messed up. As I type, the text goes below the screen and I have to go by what I think I'm writing.

OpenStudy (freckles):

I know what you mean. OS has been so annoying lately.

OpenStudy (anonymous):

so does that mean the answer should be 3/the square root of pi?

OpenStudy (mathstudent55):